Determine the first order partial derivative of the following functions:

Z=ln(x+t^2)
1
Expert's answer
2021-07-14T06:04:47-0400
"\dfrac{\partial Z}{\partial x}=\dfrac{1}{x+t^2}"

"\dfrac{\partial Z}{\partial t}=\dfrac{2t}{x+t^2}"